Circular No. 6496

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996bn IN UGC 3430 J. Mueller reports her discovery of a supernova (mag ∼ 18) on a IIIa-F

plate taken on Oct. 17 UT with the 1.2-m Oschin Schmidt telescope by K. M. Rykoski in the course of the second Palomar Sky Survey. SN 1996bn is located 2′′.7 west and 5′′.4 south of the center of UGC 3430 (α = 6^{h}15^{m}.8, δ = +64^{o}26′, equinox 2000.0). M. H. van Kerkwijk (California Institute of Technology) obtained a CCD spectrum of SN 1996bn in very poor seeing on Oct. 21 with the 5-m Hale reflector at Palomar. Preliminary inspection of the uncalibrated data by A. G. Riess, D. C. Leonard, and A. V. Filippenko (University of California, Berkeley) confirms that the object is a supernova, probably of type Ia, ∼ 1 month past maximum brightness.

V2116 OPHIUCHI R. Sood, S. James, and W. Lawson, Australian Defence Force Acade-

my; and R. Manchanda, Tata Institute, report: “Regular monitoring of the Hα emission profile of V2116 Oph, the optical counterpart of the symbiotic x-ray pulsar binary GX 1+4, has been in progress in order to establish its orbital period. Observations made with the 1.9-m telescope (+ coudé spectrograph + CCD) at Mount Stromlo Observatory on Oct. 16 failed to detect the Hα line, the first such non-detection since observations were started in 1991 Sept. Previously the spectrum had been dominated by the Hα emission, with an integrated line flux of ∼ 2.5 × 10^{−13} erg cm^{−2} s^{−1}, ∼ 100 times stronger than the continuum. Non-detection was confirmed by C. Heisler (Mount Stromlo) with the 2.3-m telescope (+ DBS spectrograph + CCD) at Siding Spring on Oct. 19, suggesting that the star was fainter than V ∼ 18. The weakening of the Hα emission coincides with the weakening of x-ray pulsations (IAUC 6478, 6488). Our previous observations made with the 1.9-m telescope on Aug. 9 detected the Hα emission at a flaring level of ∼ 10^{−12} erg cm^{−2} s^{−1}, a factor of ∼ 4 above the ambient Hα line flux, consistent with the increase in the phase-averaged pulsed x-ray flux reported on IAUC 6478. Reports of optical pulsations from this symbiotic binary at the x-ray pulse period (IAUC 6489), and the above coincidence between the Hα line flux and x-ray pulse activity, suggest a strong link between the Hα flux and the neutron star, rather than with the accretion disk. Further observations are encouraged to monitor the current optical low state.”

1996 October 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green